Write down angular frequency expression for hydrogen like atom.
1
Expert's answer
2020-11-20T09:33:16-0500

"\omega=\sqrt{\frac{e^2}{4\times\pi\times\varepsilon\times m_e\times R^3}}" , where "e" is the charge of the electron, "\varepsilon" is the electric constant, "m_e" is the mass of the electron, "R" is the radius of the orbit.